if the area of a trapezium is 28 m*m and two parallel sides are 8 m and60 dm respectively, find the altitude?
Answer by mananth(16946)

Area of trapezium = 1/2 * sum of parallel sides * height
=1/2 (8+6)*h
28=1/2 * 14 * h
28/7 = h
h= 4 m